### v1

# Further constraining the neutron star-black hole merger rate

Abstract
Current template-based gravitational-wave searches for compact binary mergers [1, 2, 3, 4] neglect the general relativistic phenomenon of spin-induced orbital precession [5]. Owing to their asymmetric masses, gravitational-waves from neutron star-black hole (NSBH) binaries are prime candidates for displaying strong imprints of spin-precession. As a result, current searches may be missing a significant fraction of the astrophysical population, and the detected NSBH population may be significantly suppressed or biased. Here we report the most sensitive search for NSBH binaries to date by including spin-precession for the first time [6]. We analyze data from the entirety of the third LIGO–Virgo–KAGRA gravitational-wave observing run [7] and show that when accounting for spin-precession, our search is up to 100% more sensitive than the search techniques currently adopted by the LIGO–Virgo–KAGRA collaboration (for systems with strong precessional effects). This allows us to more tightly constrain the rate of NSBH mergers in the local Universe. Firstly, we focus on a precessing subpopulation of NSBH mergers; the lack of observed candidates allows us to place an upper limit on the merger rate of $R_{90}=79\,\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$ with 90% confidence. Secondly, we tighten the overall rate of NSBH mergers; we show that if there is no preferred direction of component spin, the rate of NSBH mergers is on average 16% smaller than previously believed. Finally, we report four new subthreshold NSBH candidates, all with strong imprints of spin precession, but note that these are most likely to be of terrestrial origin.
{affiliations}
Institute of Cosmology and Gravitation, University of Portsmouth, Portsmouth, PO1 3FX, UK
1 Searching for neutron star-black hole mergers
Our ability to infer the neutron star-black hole (NSBH) merger rate from gravitational-wave (GW) observations is dependent on the search algorithms being used. A search algorithm that is not optimised for the detection of GWs from NSBH binaries will not identify as many GWs, and may infer an inflated merger rate with superfluous uncertainties. On the other hand, a more sensitive search may not only be able to uncover more GWs, but also explore a greater spacetime volume $\langle VT\rangle$.
The standard approach to observe GW signals from NSBH binaries is through template-based matched filter searches [1, 2, 3, 4], a well established signal processing technique. Here, a GW signal masked by noise is identified by matching the data with a pregenerated bank of templates. A limitation of template-based matched filter searches is that signals may remain undetected if they are not sufficiently similar to at least one template in the bank. For GW astronomy, templates are constructed from theoretical GW models, and to date LIGO–Virgo–KAGRA (LVK) collaboration analyses are restricted to templates that neglect the general-relativistic phenomenon of spin-induced orbital precession [5]. Spin-precession arises when there is a non-zero effective perpendicular spin [8], and hence a spin angular momentum that is mis-aligned with the orbital angular momentum of the binary. A binary with significant spin-precession has effective perpendicular spin close to unity, while a binary with spins aligned with the orbital angular momentum has effective perpendicular spin close to zero.
Restricting templates to ignore spin-precession implies that a significant fraction of the NSBH population may remain undetected [9, 10]. If we assume that current template banks can only detect NSBH binaries with minimal spin-precession (effective perpendicular spin $<0.1$), we estimate that we could be missing as many as $\sim 85\%$ of GWs from NSBH binaries formed in the local Universe (assuming an agnostic population of NSBHs, with the neutron star’s spin $<0.05$ inline with observations from Galactic binaries [11]).
Current search pipelines are more likely to miss GWs from binaries formed through the dynamic channel – binaries formed in a dense stellar environment such as globular clusters – since they are believed to have an isotropic distribution of spins and spin-precession is more likely [12]. On the other hand, current searches are tailored to detect GWs from binaries formed through the isolated channel as minimal spin-precession is expected: binaries formed through the isolated channel evolve from a pair of massive binary stars, which shed angular momentum through stable and unstable mass transfer. Although natal kicks generated through core collapse can cause small mis-alignments, we expect the spins of isolated binaries to be approximately aligned with the orbital angular momentum [13]. It is of great astrophysical interest to be able to identify, or equally rule out, NSBHs with spin-precession as it will allow the formation mechanism of NSBHs in the Universe to be determined.
In this work, we present for the first time a dedicated template-based matched filter search for NSBHs which is sensitive to the whole NSBH population (see Methods for details on our template bank and the generation process). By restricting attention to data from the third GW observing run [7], O3, we find that our search is on average $44\%$ more sensitive than the search techniques currently adopted by the LVK, with sensitivity most improved (up to 100%) for systems demonstrating strong imprints of spin-precession. This enables us to place tighter constraints on the rate of NSBH mergers in the local Universe. Firstly, we report the merger rate of a precessing subpopulation of NSBH mergers. We then further constrain the overall rate of NSBH mergers to be on average 16% smaller than the rate we compute assuming a search that requires no misalignment between component spin and orbital angular momentum (similar to those currently adopted by the LVK). Being able to observe and correctly quantify the NSBH merger rate is of significant astrophysical interest as it will allow their underlying source parameters to be measured more accurately. This enables (for example) a better understanding of the impact of supernova kicks and a much more precise measurement of the Hubble constant [14, 15, 16].
Figure 1: Percentage change in the merger rate of NSBHs in the local Universe, and its implications. Percentage change in the search sensitivity to NSBH mergers (Left) and percentage change in the merger rate of NSBHs in the local Universe (Upper Right) between a template bank which includes spin precession (generic-spin) and a template bank which restricts spins to be aligned with the orbital angular momentum (aligned-spin). The change in sensitive volume is shown as a function of the binaries total mass and the degree of spin-precession (encoded by the effective perpendicular spin $\chi_{\mathrm{p}}$). The change in NSBH merger rate is shown just as a function of the effective perpendicular spin. Assuming the full parameter space, $\chi_{\mathrm{p}}\leq 1$, we infer the NSBH merger rate to be $16\%$ smaller than previously published results. We also show the fraction of precessing NSBH binaries in the Universe as a function of $\chi_{\mathrm{p}}$ (Lower Right). According to our generic spin search, we estimate that no more than $30\%$ of NSBHs in the Universe undergo spin-precession for an independent subpopulation of NSBH binaries with $\chi_{\mathrm{p}}>0.4$.
2 Estimating the neutron star-black hole merger rate
In this work, we focus on constraining the NSBH merger rate; the NSBH merger rate is estimated through GW observations by combining the search sensitivity with the number of confident GW signals detected. To access a search’s sensitivity, a search is performed on simulated signals drawn from a reference population model that have been added to real GW detector noise. Based on the number of signals injected and recovered, a sensitivity estimate can be calculated [17]. Our results show that a dedicated search that includes spin precession is on average $44\%$ more sensitive to NSBH mergers than the search techniques currently adopted by the LVK [1]. This increases to an average of $85\%$ for highly precessing NSBH systems. This can be seen in the left-hand panel of Figure 1, which compares the sensitivity of our generic spin search with a comparable search using only aligned-spin waveforms. For highly precessing (effective perpendicular spins $>0.75$), and high total mass NSBH systems (total mass $>14\,M_{\odot}$), we observe $>100\%$ increase in the sensitive volume. Although it is possible to marginalize over multiple discrete population models when estimating the sensitive volume, we assume a single, agnostic population owing to a large uncertainty in the underlying NSBH population (see Methods). This analysis ignores the possibility of the neutron star being tidally disrupted during merger and deformed during the inspiral. Tidal disruption is unlikely to play a pivotal role, since most NSBH binaries are likely to merge before the neutron star is disrupted for any expected neutron star equation of state [18]. Likewise, searches that ignore tidal deformability are likely to capture more than 99% of GW signals from NSBH binaries [19].
Figure 1 shows the largest difference in sensitive volume for NSBH systems with large effective perpendicular spin. Although the maximum neutron star spin observed from Galactic binary neutron star mergers is $<0.05$ [11] we argue that highly precessing NSBH binaries are astrophysically possible. Given that the effective perpendicular spin is highly dependent on the dynamics of the primary object for asymmetric component mass binaries, highly precessing NSBH systems require the black hole to have large spin magnitudes that are misaligned with the orbital angular momentum [8]. Firstly, highly spinning black holes are possible: although the current black hole population estimates from GW observations point towards low spin magnitudes, with half of spin magnitudes less than 0.26 (based on binary black hole mergers) [20], rapidly spinning black holes have been observed through GW [21, 22] and X-ray observations [23, 24, 25]. In fact, Cygnus X-1 is expected to form an NSBH binary with a near-maximal spin black hole [26]. Secondly, spin magnitudes misaligned with the orbital angular momentum are possible: NSBH binaries formed through dynamic capture are expected to produce precessing binaries [12]. Similarly, although binaries formed in isolation are expected to have spins primarily aligned with the orbital angular momentum [13], precessing binaries are possible in certain regions of the parameter space [27], and as a consequence of supernova kicks [12, 28] and subsequent evolution via mass transfer [29].
The LVK employs a search that restricts spins to be aligned with the orbital angular momentum, and covers binaries with masses between $2\,M_{\odot}$ and $500\,M_{\odot}$ [22]. We generated a reduced version of this LVK search by similarly restricting spins to be aligned with the orbital angular momentum, but covering the same NSBH parameter space as our generic spin search. While the full LVK search is also configured to find GW signals observed in one [30] or more [1] GW detectors, we limit all searches in this study to find coincident signals in two GW detectors. Since they are the most sensitive, we restrict attention to data collected by LIGO-Hanford and LIGO-Livingston [31].
Once the search sensitivity has been estimated, the number of confident GW signals found by our search must be determined. A confident GW signal is an event of astrophysical origin with a false alarm rate smaller than $10^{-2}\,\mathrm{yr}^{-1}$. The false alarm rate is defined as how often an event of equivalent or greater significance would occur. To date, the LVK has reported several GW signals from NSBH binaries [32, 22, 33]. However, our search found only one confident detection: (see Figure 3 in the Supplementary material): GW200115_042309 [33], a multi-detector NSBH candidate reported by the LVK collaboration [33, 22, 34]. We did not find GW200105_162426 [33] since this was a single detector candidate when ignoring the data collected by Virgo [35]. Our search found GW200115_042309 with a FAR of $1.8\times 10^{-5}\,\mathrm{yr}^{-1}$ and a probability of astrophysical origin $99.9\%$. Through Bayesian inference techniques, GW200115_042309 was inferred to have minimal precession; it was found that the effective perpendicular spin $\lesssim 0.4$ [33].
Once the search sensitivity has been estimated and the number of confident GW observations determined, the merger rate density can be calculated. Given that our generic-spin search did not find any evidence for NSBH binaries with effective perpendicular spin $>0.4$, we place upper limits on the merger rate of a theoretical subpopulation of precessing NSBHs in the local Universe, see Methods for details. Defining a precessing subpopulation of NSBH binaries as effective perpendicular spin $>0.4$, we infer an upper limit on the merger rate at 90% confidence to be $R_{90}=79\,\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$. This reduces to $R_{50}=13\,\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$ at 50% confidence. Our generic-spin search reduces the upper limit on the merger rate by 40% compared to a comparable search using only aligned-spin waveforms.
Our generic-spin search also allows us to place tighter constrains on the overall rate of NSBH mergers in the local Universe. Assuming a Poisson likelihood over the astrophysical rate, a Jeffreys prior, and marginalizing over a 15% uncertainty arising from calibration uncertainties from the GW detectors, we tighten the merger rate density of NSBHs in the local Universe to $44^{+102}_{-37}\,\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$; a reduction of $16\%$ on average compared to the merger rate estimated when using a search with no spin-precession (similar to those currently adopted by the LVK) [1] $52^{+130}_{-44}\,\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$. Although our reported merger rate density assumes a single confident detection and data restricted to O3 only, we expect a similar reduction when more confident detections and data are included. Indeed, if we assume two confident detections, we similarly observe a 16% reduction in the inferred NSBH merger rate.
In the upper right-hand panel of Figure 1, we show the change in the NSBH rate as a function of the effective perpendicular spin. Assuming negligible spin-precession (effective perpendicular spin $\sim 0$) we find comparable merger rate estimates between our search and an equivalent LVK analysis. This is expected since the change in sensitive volume for NSBHs with negligible spin-precession remains close to 0. However, as soon as spin-precession becomes significant (effective perpendicular spin $>0$), we infer a smaller merger rate than previously believed. When considering the fully precessing parameter space, we observe a merger rate that is 16% lower than that estimated with the tools currently adopted by the LVK. We report that the change in the NSBH merger rate is approximately linear in the effective perpendicular spin.
Name GPS time False Alarm Rate SNR $p_{\mathrm{astro}}$ (%)
GW190421_203205 1239913943.259 $6.7\,\mathrm{yr^{-1}}$ 9.7 0.7
GW190916_134302 1252676600.135 $10.0\,\mathrm{yr^{-1}}$ 9.5 0.5
GW191225_091958 1261300816.100 $8.3\,\mathrm{yr^{-1}}$ 9.8 0.5
GW2003010_030555 1267844773.573 $10.0\,\mathrm{yr^{-1}}$ 9.7 0.4
Table 1: Properties of additional sub-threshold NSBH candidates Properties of the sub-threshold NSBH candidates uniquely found by our search. For each candidate we report the GPS time, False Alarm Rate, signal-to-noise ratio (SNR), and probability of astrophysical origin $p_{\mathrm{astro}}$.
3 Impact on the underlying NSBH population
Quantifying the fraction of precessing NSBH binaries in the Universe has significant implications for understanding the formation mechanism of NSBHs. We estimate an upper limit on the fraction of precessing NSBH binaries by directly comparing the rate of precessing binaries (assuming zero observations) with the expected rate of NSBH mergers for the population as a whole. If the two rate estimates are in agreement, zero precessing NSBH binaries are expected in the Universe. In the bottom right-hand panel of Figure. 1, we show the fraction of precessing NSBH binaries as a function of $\chi_{p}$. Assuming an independent subpopulation of precessing NSBH binaries that has $\chi_{p}>0.4$, we estimate that no more than 30% of NSBH binaries in the Universe undergo spin-precession. This increases to $35\%$ for $\chi_{\mathrm{p}}>0.8$. Owing to the improved sensitivity of our generic-spin search, we see a significant reduction in the estimated fraction of precessing NSBH binaries compared to that obtained fom an aligned-spin search (similar to those currently adopted by the LVK). Assuming an independent subpopulation of precessing NSBH binaries with $\chi_{p}>0.8$, ignoring all prior beliefs, we reduce the expected fraction of precessing NSBH binaries in the Universe by $\sim 40\%$. We note that this is an indicative measurement and does not account for the uncertainties on the rate estimates.
Inferring the underlying population properties of NSBH binaries in the Universe requires a careful understanding of the selection biases from search algorithms. The reason is because our searches are more likely to observe certain binary configurations than others, meaning that the number of observations in a specific part of the parameter space is not necessarily representative of the underlying population; for example, searches are more likely to observe binaries with spins aligned with the orbital angular momentum than binaries with spins anti-aligned with the orbital angular momentum. Given that our dedicated generic-spin search is on average 44% more sensitive to NSBH mergers than the equivalent LVK analysis, see Figure 1, we investigate how this impacts the underlying NSBH population inferred through hierarchical Bayesian inference. The details of our Bayesian analysis are given in Methods.
When employing a powerlaw distribution for the mass ratio [36], we find no qualitative difference between the inferred population properties when using a more sensitive search; although a marginally larger Bayes factor is obtained when using the generic-spin search results. This is expected since existing results account for the search sensitivity, and a less sensitive search should only give larger uncertainties reflecting the lack of information in a given region. It is possible that larger differences will be seen with more observations, especially if our generic-spin search uncovers additional NSBH observations that are not found with the search pipelines currently employed by the LVK.
4 Marginal NSBH observations
We report four new sub-threshold NSBH candidates, GW190421_203205, GW190916_134302, GW191225_091958, GW2003010_030555, all with false alarm rates $10\,\mathrm{yr}^{-1}<\mathrm{FAR}<6\,\mathrm{yr}^{-1}$ and probability of astrophysical origin $p_{\mathrm{astro}}<1\%$. Each candidate’s FAR, signal-to-noise ratio and probability of astrophysical origin is given in Table 1. Given the small probability of astrophysical origin, it is likely that these are of terrestrial origin and were therefore not included in our merger rate estimate.
Since our generic-spin search found these marginal candidates, it is likely that they exhibit strong evidence for spin-precession. If real, they are astrophysically significant as they differ from current population estimates for NSBHs [36], which favours the isolated formation mechanism [37]. We perform Bayesian inference, under the assumption that these sub-threshold NSBH candidates are of astrophysical origin, to infer the properties of each source. The details of our Bayesian analyses are given in Methods.
Figure 2 shows the posterior probability distribution for three of the four candidate’s primary mass and effective perpendicular spin. We do not include GW190421_203205 since its posteriors are highly peaked, indicating a possible issue during sampling. This makes it difficult to inspect the 90% credible interval; we infer the primary mass to be $8.1^{+0.7}_{-0.2}\,M_{\odot}$ and effective perpendicular spin $0.228^{+0.003}_{-0.003}$. We find that all but one candidate has significant spin-precession, with effective perpendicular spin constrained away from 0, and away from the prior indicating an informative measurement. If real, these newly reported marginal NSBH candidates indicate a distinct population likely formed through dynamic capture.
5 Conclusions
Figure 2: The primary mass and effective perpendicular spin of 3 new sub-threshold NSBH candidates. Corner plot showing the marginalized posterior probability for the primary mass and effective perpendicular spin for the newly reported sub-threshold NSBH candidates. Contours represent the 90% credible interval. The prior for the effective perpendicular spin is shown for comparison.
In this work, we present for the first time a dedicated template-based matched filter search for NSBHs that is sensitive to the whole NSBH population. We apply our search to all data collected during the third GW observing run, O3, and place upper limits on the merger rate of a theorised subpopulation of precessing NSBH candidates. We further constrain the overall NSBH merger rate, and show that on average the inferred NSBH merger is reduced by 16% compared to the results we compute when assuming a search that requires no misalignment between component spin and orbital angular momentum (similar to those currently adopted by the LVK). Although we only considered data collected during O3, we envisage a similar reduction in merger rate when more data are considered. We further identify four new subthreshold NSBH candidates, all with strong imprints of spin-precession. However, we note that these are most likely to be of terrestrial origin.
Given that our more refined NSBH search contains only 3$\sim$ 3$\times$ 3 × more templates than a reduced search currently adopted by the LVK, we highly encourage its use in the future. Our search has the potential to determine the formation mechanism of NSBHs in the local Universe, as well as constrain properties of the surrounding environment. For instance, NSBHs formed in disks of active galactic nuclei (AGN) is a promising formation channel since its predicted merger rate is consistent with observations: $10-300\,\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$ [38]. Tighter constrains on the NSBH merger rate from our search will further restrict models for the mass function, mass segregation and populations in AGN disks. Moreover, our generic-spin search may also discover precessing NSBHs that would otherwise remain undetected. Confirmation that NSBHs form through the dynamic channel would decrease the probability that NSBHs form in young star clusters, another promising formation channel, since it is expected that most systems forming coalescing NSBHs are ejected without undergoing dynamic exchanges, and hence have minimal spin-precession [39].
6 Methods
6.1 Matched filter searches:
Template-based matched filter searches identify gravitational-wave (GW) signals by matching the data against “template” waveforms representing predicted gravitational-wave signals. A signal-to-noise ratio (SNR) [40, 41] can be computed by matched filtering the data against a given template waveform, $h_{i}$ according to
$\rho^{2}=\frac{(h_{i}|s)^{2}}{(h_{i}|h_{i})}.$ (1)
Here $s$ is the data and $(a|b)$ is the complex noise-weighted inner product timeseries, defined as
[…]
where $S_{n}(f)$ is the power spectral density (PSD). If one was to know exactly the signal, $h_{i}$, that is being searched for then the SNR, added in quadrature, directly relates to the probability that the data actually contains that signal. An SNR of 5, for example, is interpreted as a 5-sigma confidence in the signal being present under the assumption that the noise is Gaussian and stationary.
However, a significant issue for these matched-filter searches is the dimensionality of the problem. A binary black hole merger, on a circular orbit, is described by fifteen parameters, and one must be able to observe the signal anywhere in this fifteen dimensional parameter space. In current LVK searches, the dimensionality of the problem is decreased by assuming that the angular momentum of the two component bodies is aligned, or anti aligned, with the orbit. One is then able to analytically maximise the SNR over the binary’s sky location, orientation, luminosity distance, and time of coalescence, reducing the problem to 4 dimensions; the two component masses, and two component spin magnitudes.
A bank of template waveforms is then produced across this 4-dimensional parameter space, such that for any signal at least a certain fraction of the total signal power available is recovered, normally chosen to be at least 97% of the maximum possible SNR. One then filters the bank of waveforms against the available data, identifying times where the SNR is elevated. It is important not to assume that the data is either Gaussian or stationary and therefore a set of signal-based consistency tests are applied to attempt to reject, or downweight, times where the elevated SNR is due to instrumental artefacts, or “glitches”. Finally, a background rate of triggers is empirically estimated from the data and potential events are assigned a significance, representing our belief in whether or not they are real. An interested reader can find a much more detailed description of current methods in the relevant methods papers [40, 41, 1, 2, 42, 3, 4].
While undeniably successful, this technique does result in a significant loss in sensitivity for signals where spins are misaligned with the orbit, in particular systems where this misalignment causes a precession of the orbital plane, which is not modelled by the waveform filters [10]. We recently proposed a novel method for incorporating misaligned spins in the search [6] and we apply that method here to carry out a search for NSBH mergers with misaligned spins in data from the third LVK observing run. We briefly summarize the method here for the clarity of the reader, but point the reader to [6] for the complete description.
When searching for signals that include spin-precession we must first construct a bank of templates. The standard technique to maximize over orientation and sky location parameters can no longer be used since the extrinsic parameters affect the morphology of a precessing GW signal. One can avoid this problem by building a template bank over a parameter space with significantly increased dimension, but this significantly increases the computational cost of the search, proving impractical for NSBH systems where the impact of precession is particularly important [10, 43]. It was recently demonstrated that precessing signals can be constructed by utilizing a “harmonic decomposition” of a gravitational wave signal [44]. Here, a gravitational wave signal can be written as a power series of non-precessing templates, meaning that the standard maximisation routine can be applied to each harmonic individually [6]. By only filtering over the most important 2, or 3, harmonics of a precessing template the template bank can be kept small, while still retaining sufficient sensitivity to observe precessing systems [6]. We use this technique to produce a template bank which is only 3$\sim$ 3$\times$ 3 × larger than aligned spin template banks. We then implement the harmonic decomposition into the matched filtering process, by separately filtering each of the harmonics for each signal, before recombining and computing significance. We have adapted the signal-based consistency tests, and the method by which significance is computed to account for the harmonic decomposition [6]. In our method paper, we argued that this method would increase the sensitive volume increased by $\sim 60\%$ on average across the full generic spin parameter space, and $\sim 100\%$ for strongly precessing GW signals [6].
In this work, we apply our generic spin search technique to perform the most sensitive search for neutron-star black hole signals to date. We searched for signals in data from the third LVK gravitational-wave observing run (made available through the Gravitational-Wave Open Science Center, GWOSC). We search through the entirety of this data twice: First, using a template bank with spins aligned with the orbital angular momentum to set a baseline, i.e. the current technique adopted by the LVK, and second, using a precessing template bank. We assess the increased performance by comparing each search’s sensitivity to NSBH mergers. In order to evaluate a search’s sensitivity, simulated signals are added to the data and searched for in the same way as real signals. In this work, we draw the primary mass in the detector frame uniformly between $5\leq m_{1}\leq 20.0$ and the secondary mass uniformly between $1.2\leq m_{2}\leq 1.7$, while ensuring that the total mass $6.2\leq M\leq 21.7$. The spin magnitudes are drawn uniformly between 00 and $0.99$ for both objects, and the spin orientations, sky location and orientation angles are drawn isotropically. The signals are generated with the IMRPhenomXP waveform model [45].
IMRPhenomXP models precessing GW signals by performing a time-dependent rotation to an aligned-spin waveform [46] to mimic precession effects [47]. While the accuracy of IMRPhenomXP generally worsens as the component masses become more asymmetric, when compared against numerical simulations of General Relativity [48], we argue that this will have a minimal effect on our search. The reason is because templates in our bank are chosen to guarantee that GW signals are identified up to a given threshold. This threshold is larger than the average accuracy of IMRPhenomXP [48].
We consider a simulated signal found if the false-alarm rate is less than 1 per 100 years; the false-alarm rate is a commonly used quantity by the LVK to assess the significance of detected GW signals. We compute the detection efficiency using fifty distance bins, and the volume contained in each distance bin is then multiplied by its detection efficiency. The results are then summed across all bins to calculate the sensitive volume. We compute the change in sensitive volume by comparing the sensitivity from the first and second searches, i.e. a search that uses a template bank with spins aligned with the orbital angular momentum, and a search using a precessing template bank.
The change in sensitive volume as a function of the total mass and in-plane spin of the binary is shown in Figure 1.
Figure 3: Cumulative histogram of GW triggers. Comparison between expected background (dashed black line) and coincident foreground GW triggers (black stars). The shaded regions represent $1\sigma$, $2\sigma$, $3\sigma$ and $4\sigma$ counting errors.
6.2 Inferring the merger rate:
A coincident trigger occurs when more than one GW detector observes the same candidate within a time frame of a few milliseconds or less (with the exact timing dependent on the GW network considered), and it can be classified into one of two categories: of astrophysical origin or those from terrestrial sources such as instrumental or environmental effects. A coincident trigger of astrophysical origin is assumed to follow a Poisson process with mean $\zeta=R\langle VT\rangle$, where $R$ is the rate binary mergers and $\langle VT\rangle$ is the search sensitivity. Assuming all coincident triggers have zero probability of terrestrial origin, and a Poisson likelihood, a probability distribution for the mean number of triggers $n$ for a given $\zeta$ can be obtained with
$p(n|\zeta)=\frac{\zeta^{n}e^{-\zeta}}{n!}.$ (3)
Through Bayes’ theorem, this can be re-expressed to calculate a probability distribution for $\zeta$,
$\begin{split}p(\zeta|n)&\propto p(n|\zeta)\,p(\zeta)\\ &\propto\frac{\zeta^{n}e^{-\zeta}}{n!}\,p(\zeta)\\ \end{split}$ (4)
where $p(\zeta)$ is the prior for $\zeta$. Equation 4 can be rewritten as a joint probability on the astrophysical rate and sensitive volume,
$p(R,\langle VT\rangle|n)\propto\frac{(R\langle VT\rangle)^{n}e^{-R\langle VT\rangle}}{n!}\,p(R,\langle VT\rangle),$ (5)
which can be simplified by noting that $p(R,\langle VT\rangle)=p(R|\langle VT\rangle)\,p(\langle VT\rangle)$, where $p(\langle VT\rangle)$ is the prior for $\langle VT\rangle$. Although other priors are possible, we impose a Jeffreys prior in this work for $p(R|\langle VT\rangle)$: $\propto(R\langle VT\rangle)^{-1/2}$. We note that the choice of prior will likely have a non-trivial effect on the inferred merger rate for a small number of confident triggers. Equation 5 then takes the form,
$p(R,\langle VT\rangle|n)\propto p(\langle VT\rangle)\,\frac{(R\langle VT\rangle)^{n}e^{-R\langle VT\rangle}}{n!\sqrt{R\langle VT\rangle}}$ (6)
from which a probability distribution for the merger rate can be inferred by marginalizing over $\langle VT\rangle$,
$p(R|n)=\int{p(R,\langle VT\rangle|n)\,d\langle VT\rangle}$ (7)
Up until now, the GW detectors have been assumed to be perfect instruments. However, GW detectors are subject to an uncertainty in their calibration, leading to an uncertainty in our estimate of the search sensitivity [49]. This uncertainty can be incorporated into our analysis by assuming a log normal prior for $\langle VT\rangle$,
$p(\log\langle VT\rangle)\propto\mathcal{N}\left(\log\mu,\frac{\sigma}{\mu}\right)$ (8)
where $\mu$ is our estimate for the search sensitivity and
$\sigma^{2}=\sigma_{\mathrm{cal}}^{2}+\sigma_{\mathrm{stat}}^{2}$ (9)
where $\sigma_{\mathrm{cal}}$ is the uncertainty in $\langle VT\rangle$ arising from the calibration of GW detectors, and $\sigma_{\mathrm{stat}}$ is the estimated uncertainty from searching for a finite number of simulated signals that have been added to the GW data. In this work, we assume $\sigma_{\mathrm{cal}}=0.15\mu$, i.e. a 15% uncertainty in our estimated volume.
When no confident GW signals are observed, an upper limit on the rate with confidence $c$ can be obtained by solving,
[…]
When more than one confidence GW signal is observed, an estimate for the merger rate can be obtained with,
$p(R|n=1)\propto\int\sqrt{R\langle VT\rangle)}e^{-R\langle VT\rangle}\,\mathcal{N}\left(\log\mu,\frac{\sigma}{\mu}\right)d\langle VT\rangle$ (11)
which can be inferred through Markov Chain Monte Carlo sampling [50]. In this work we searched for GWs from NSBH binaries in O3111O3 was divided into two halves where the first half (O3a) ran between 1 April 2019, 1500 UTC and 1 October 2019, 1500 UTC, while the second half (O3b) ran between 1 November 2019, 15:00 UTC and 27 March 2020, 17:00 UTC , and found a single confident GW signal. Figure 3 compares the observed distribution of events to the background distribution used to assess significance. We see that only one coincident trigger is more significant than any background event in the data, with significance estimated to be greater than $4\sigma$.
6.3 Inferring the source properties of marginal candidates:
Assuming a GW candidate is of astrophysical origin, Bayesian inference can be performed to infer the properties of the source $\bm{\lambda}=\{\lambda_{1},\lambda_{2},...,\lambda_{16}\}$; $\bm{\lambda}$ is a sixteen dimensional vector for a NSBH binary on a circular orbit. The additional parameter beyond a quasi-circular binary black hole is the tidal deformability of the secondary object. Bayesian inference involves calculating the model dependent posterior probability distribution function, defined as
$p(\bm{\lambda}|s,\mathfrak{M})=\frac{\mathcal{L}(s|\bm{\lambda},\mathfrak{M})\,\Pi(\bm{\lambda}|\mathfrak{M})}{\mathcal{Z}},$ (12)
[…]
$\mathcal{L}(s|\bm{\lambda},\mathfrak{M})=\exp\left(-\frac{1}{2}(s-\mathfrak{M}(\bm{\lambda})|s-\mathfrak{M}(\bm{\lambda}))\right)$ (13)
where $(a|b)$ is defined in Equation 2.
Unfortunately, it is often not possible to analytically calculate the posterior probability distribution function for GW astronomy. The reason is because the evidence involves evaluating a sixteen dimensional integral. For this reason, stochastic sampling techniques such as Nested sampling [52, 53] or Markov-Chain Monte-Carlo [50], are often employed. Here, the sixteen dimensional parameter space is explored by evaluating the likelihood in randomly chosen points in the parameter space, and eventually the most probable region of parameter space is found through an iterative process.
In this work, we estimate the source properties of marginal NSBH candidates by evaluating Equation 12 through Nested sampling. We employ the dynesty nested sampler [54] via bilby [55, 56] as has been done in all LVK analyses since the third GW catalog [22]. We employ agnostic priors for all parameters, identical to those used in all detections made by the LVK [22] – uniform in the component masses, spin magnitudes and cosine of the spin tilt angles – and assume 2000 live points with the bilby-implemented rwalk sampling algorithm with an average of 60 accepted steps per MCMC for all analyses. We analyse 64s of data including frequencies in the range of 20–896Hz in the calculation of the likelihood. To reduce the computational cost we employ a multi-banded likelihood [57], which takes advantage of the chirping behaviour of GW signals to reduce the number of time samples at which waveforms are evaluated. The multi-banded likelihood has been shown to give small log likelihood ratio differences, even for waveform models that include precession and higher order multipole content [57].
To calculate the posterior distribution, a GW model $\mathfrak{M}$ must be chosen and the PSD $S_{n}(f)$ estimated. In this work, we employ one of the most accurate and cutting edge models currently available: IMRPhenomXPHM [45]. Although this model is designed for GWs from binary black holes, we use it to estimate the source properties of possible NSBH candidates since current NSBH models do not include spin-precession [58, 59]. This is also the model which has been used in recent LVK analyses of potential NSBH binaries observed in the fourth observing run [34]. As such, we perform Bayesian inference on a fifteen dimensional parameter space, and neglect the tidal deformability of the secondary object. This is unlikely to have significant effect on Bayesian inference analyses [60].
To remain consistent with LVK analyses, we use BayesWave [61] to estimate the on-source PSD for data surrounding each marginal candidate. We also marginalize over calibration error by using publicly available envelopes [62]. Envelopes are not available for all times in O3, so we use those that have been released alongside GW signals [63] occurring at a similar time to the events reported in Table 1.
6.4 Inferring the properties of the underlying population:
Assuming a set of measurements for $N_{\mathrm{obs}}$ GW observations, hierarchical Bayesian inference can be used to infer the properties of the underlying population $\bm{\Lambda}=\{\Lambda_{1},\Lambda_{2},...,\Lambda_{n}\}$. A posterior distribution for the hyper parameters of the population model can be inferred through Bayes’ theorem via,
$p(\bm{\Lambda}|\{D\})=\frac{\mathcal{L}(\{D\}|\bm{\Lambda})\,\Pi(\bm{\Lambda})}{\mathfrak{Z}}$ (14)
[…]
where $D_{i}$ is the $i$th event in the catalog, and $p_{\mathrm{det}}$ is the detection probability of a GW with parameters ${\bm{\lambda}}_{i}$.
The detection probability is directly related to the performance of the search; a search that is not sensitive to spin-precession will have a small detection probability for a strongly precessing NSBH, while a generic-spin search will have a larger detection probability. The detection probability is often estimated by re-using the search results from a randomly chosen set of simulated signals drawn from a population model with hyper parameters ${\bm{\Lambda}}_{\mathrm{ref}}$ that was used to estimate the searches sensitive volume [65, 17].
In this work we estimate the properties of the underlying NSBH population through Nested sampling. We employ the Dynesty [54] nested sampler via GWPopulation [66] and Bilby [55] as has been done in recent LVK analyses. We use the same priors and population models as in Biscoveanu et al. [36]. We assume 1000 live points and consider a catalog containing GW200105_162426 and GW200115_042309 [33]. We do not include other NSBH detections made by the LVK since it has been shown that the underlying population properties remain comparable when only these two highly significant events are included [36]. Despite our generic-spin search not being able to isolate GW200105_162426, we include it in this analysis since it remains influential in the hierarchical Bayesian analysis. We use the publicly available mixed posterior samples for GW200105_162426 and GW200115_042309 [63].
Acknowledgements
We thank Gareth Cabourn Davies, Thomas Dent, Rahul Dhurkunde and Francesco Pannarle for comments on this manuscript, and Laura Nuttall for discussion throughout this project. C.H thanks Zoheyr Doctor, Vicky Kalogera and the Center for Interdisciplinary Exploration and Research in Astrophysics (CIERA) for hosting C.H in 2023, and for their guidance on running GWPopulation. I.H thanks the STFC for support through the grants ST/T000333/1 and ST/V005715/1, and C.H thanks the UKRI Future Leaders Fellowship for support through the grant MR/T01881X/1. We are also grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459, as well as the SCIAMA High Performance Compute (HPC) cluster which is supported by the ICG and the University of Portsmouth. This material is based upon work supported by NSF’s LIGO Laboratory, which is a major facility fully funded by the National Science Foundation. This research has made use of data, software and/or web tools obtained from the Gravitational Wave Open Science Center (https://www.gw-openscience.org), a service of LIGO Laboratory, the LIGO Scientific Collaboration and the Virgo Collaboration. LIGO is funded by the U.S. National Science Foundation. Virgo is funded by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale della Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by Polish and Hungarian institutes.
Data Availability
All GW data analysed in this work can be obtained through the Gravitational Wave Open Science Center (https://www.gw-openscience.org). Details regarding our generic-spin NSBH search code, including template bank generation and running our search of a chunk of O3 data, can be found at: https://icg-gravwaves.github.io/precessing_search_paper/. Posterior samples obtained for the marginal NSBH observations detailed in Table. 1 can be found at: https://icg-gravwaves.github.io/nsbh_search_on_O3_paper/.
Author contributions
I.H. initiated and lead the study. I.H performed the search on O3 data with a template bank constructed from templates which include spin-precession, and C.H performed the search on O3 data with a template bank which neglected spin-precession. C.H calculated the NSBH merger rate estimate, and performed the single-event and hierarchical Bayesian analyses on real detector data. C.H led the writing of the paper, and I.H and C.H both contributed to the interpretation of the results.
Competing interests
The authors declare that they have no competing financial interests.